Becton Labs, Inc., produces various chemical compounds for industrial use. One compound, called Fludex, is prepared using an elaborate distilling process. The company has developed standard costs for one unit of Fludex, as follows:

 


Standard Quantity
or Hours
Standard Price
or Rate
Standard Cost
Direct materials2.20ounces$23.00per ounce$50.60
Direct labor0.70hours$12.00per hour
8.40
Variable manufacturing overhead0.70hours$3.00per hour
2.10
Total standard cost per unit




$61.10

 

During November, the following activity was recorded related to the production of Fludex:

 

  1. Materials purchased, 11,000 ounces at a cost of $237,600.

  2. There was no beginning inventory of materials; however, at the end of the month, 2,650 ounces of material remained in ending inventory.

  3. The company employs 18 lab technicians to work on the production of Fludex. During November, they each worked an average of 190 hours at an average pay rate of $10.50 per hour.

  4. Variable manufacturing overhead is assigned to Fludex on the basis of direct labor-hours. Variable manufacturing overhead costs during November totaled $6,200.

  5. During November, the company produced 3,750 units of Fludex.

 

Required:

1. For direct materials:

a. Compute the price and quantity variances.

b. The materials were purchased from a new supplier who is anxious to enter into a long-term purchase contract. Would you recommend that the company sign the contract?

 

2. For direct labor:

a. Compute the rate and efficiency variances.

b. In the past, the 18 technicians employed in the production of Fludex consisted of 5 senior technicians and 13 assistants. During November, the company experimented with fewer senior technicians and more assistants in order to reduce labor costs. Would you recommend that the new labor mix be continued?

 

3. Compute the variable overhead rate and efficiency variances.
Answer #1
Answer #2
Direct materials used 8350 =11000-2650
Actual labor hours 3420 =18*190
1
a
Materials price variance 15400 F =237600-(11000*23)
Materials quantity variance 2300 U =23*(8350-3750*2.2)
b
Yes, as materials variance is favorable
2a
Labor rate variance 5130 F =3420*(10.5-12)
Labor efficiency variance 9540 U =12*(3420-3750*0.7)
b
No, as labor variance is unfavorable
3a
Variable overhead rate variance 4060 F =6200-(3420*3)
Variable overhead efficiency variance 2385 U =3*(3420-3750*0.7)